OpenStudy (anonymous):

Given F (x) = -3+4 , what is F (2) ? A. –28 B. –24 C. –2 D. 10

OpenStudy (anonymous):

will fan and medal!

OpenStudy (mindblast3r):

is this the actual question? can you post a screenshot?

OpenStudy (anonymous):

there is no pic. This is the question

OpenStudy (anonymous):

But @Mindblast3r since your good with these kind of questions could you answer some others?

OpenStudy (usukidoll):

where the heck is the x in here?

OpenStudy (usukidoll):

is it f(x) = -3x+4 or f(x) = -3+4x?

OpenStudy (anonymous):

idk.. its the question excact how i see it

geerky42 (geerky42):

Pretty sure it's \(F(x) = -3x+4\)

geerky42 (geerky42):

There is x after "-3," right? @Swordsman1305

OpenStudy (mindblast3r):

this question don't make sense lolol

OpenStudy (usukidoll):

yeah ^ but I'm not going to outright blast the answer on here

OpenStudy (anonymous):

OOOH!! YOUR RIGHT.. I DIDNT TYPE IT!!!

OpenStudy (anonymous):

its x after the 3

OpenStudy (anonymous):

sorreh

OpenStudy (usukidoll):

since f(2), x = 2... you plug x =2 into -3x+4

OpenStudy (mindblast3r):

knew it, it just proves that we got the skills to detect that it doesn't make sense

OpenStudy (anonymous):

lol

OpenStudy (usukidoll):

f(x) = -3x+4 .. since you are given f(2) f(x) => f(2) implies that x = 2 plug in 2 into the function

OpenStudy (anonymous):

so c?

OpenStudy (usukidoll):

yeah

OpenStudy (usukidoll):

f(2) = -3(2)+4 = -6+4 = -2

OpenStudy (anonymous):

thanks!!!